# Influence of catastrophes and hidden dynamical symmetries on ultrafast backscattered photoelectrons

### v1

@@ III Photoelectron momentum distributions and momentum mapping
For simplicity, in the figures that follow, we will focus on the orbits leading to the primary rescattering ridge whose energy is around $10U_{p}$, but the arguments employed in the subsequent analysis are general and can be applied to the remaining pairs of orbits. Furthermore, we have considered the initial momenta in the upper half-plane. For the orbits associated with the primary rescattering ridge, this leads to final momenta in the lower half plane, of which we just show a single quadrant (see upper panels in Figs. 4 and 6). In Fig. 3, we have used the upper half plane for the final momenta to construct the PMD, which means that the initial momenta arein the lower half-plane. However, this does not alter the subsequent conclusions as the PMDs are symmetric with regard to a reflection upon the $p_{f\parallel}$ axis. For the specific case of a monochromatic field, we also expect a reflection symmetry with regard to the $p_{f\perp}$ axis for the momentum mapping. Therefore, we have only included ionization times in a single half-cycle without loss of generality.
@@
We have observed that, for this particular pair of orbits and potential, the short (long) orbits have an odd (even) number of focal points, each of which occurs at a time when the Jacobian is vanishing. Not all focal points will cause relevant dynamical changes in the system. Some focal points occur in pairs, so that the Maslov phases associated with them will cancel each other. They are not directly related to rescattering. However, the remaining focal points will appear at the same time as a spike in the Jacobian, which is associated with rescatering and is important for the dynamics. Both orbits undergo one act of rescattering, and exhibit a single spike in the Jacobian. Physically, a spike indicates sudden changes in the electron momentum. A summary of these findings is provided in Table 2.
The corresponding initial momentum maps are given in Fig. 4(b), where the same color code as for the final momenta was used. The overall momenta for the long orbit is much smaller, which is consistent with ionization times being closer to the peak of the field. Moreover, the density of points is highest close to the polarization axis and is lowest near and at the boundary between the black and yellow regions. The high density is in line with the expected dynamics for a linearly polarized field, as ionization will happen predominantly close to the axis. Nonetheless, these regions are distorted by the presence of the Coulomb potential.
@@
Figs. 6(a)(i)-(iv) show the final momentum maps associated with each of these solutions. A noteworthy feature is that there are no solutions close to the polarization axis. This confirms that the behavior observed for the PMD in Fig. 3 is due to a change in kinematic constraints. The figure also shows that, for the soft-core case, different sheets of solutions obtained for the final momenta merge in several places, not only the rescattering ridge. The boundaries at which different solutions are glued together are represented by the dashed lines in each panel, and the color indicates to which solution it is connected. For instance, in Fig. 6(a)(i), we see that the solution $4^{(s)}_{s}$, whose final momenta are represented by the yellow scatter plot, still merges at the ridge with the long orbit $4^{(s)}_{l}$, whose final momentum mapping is given by the black scatter plot in Fig. 6(a)(ii). This behavior resembles that of their hard-core counterparts. However, in addition to that, $4^{(s)}_{s}$ is also connected to the solution $4^{(s)}_{s^{\prime}}$ [Fig. 6(a)(iii)] as the green dashed line indicates. The solution $4^{(s)}_{s^{\prime}}$, on its turn, is also connected to the $4^{(s)}_{l^{\prime}}$ solution [shown in Fig. 6(a)(iv)] via the rescattering ridge. Solutions $4^{(s)}_{l^{\prime}}$ and $4^{(s)}_{l}$ coalesce at the boundary defined near the minimal rescattering angle. All four solutions merge at a single point, where the ridges separated by half a cycle intersect. This point also marks the apex of a caustic associated with type 3 orbit [31, 43].
@@
Figure 5: Three dimensional representation of the initial perpendicular momentum map as a function of the electron’s final momentum components, calculated for the same parameters as in the previous figures, and a softening of $\alpha=1\times 10^{-6}$ [panel (a)] and $\alpha=1\times 10^{-2}$ [panels(b), (c) and (d)]. The colors of the surfaces match those used in the two-dimensional projections in Figs. 4 and 6, and the arrows indicate the merging of two or more sheets.
@@
This merging is better seen in the full three-dimensional surface, displayed in Figs. 5(b) to (d). The arrows in Fig. 5(b), indicate the coalescence of two solutions along the fold introduced by the softening. The left arrow shows the merging of the red and black branches of solutions, closest to the $p_{f,\parallel}$ axes, while the right arrow indicates the fusion of the yellow and green branches. The coalescence of these sheets is determined by the minimal rescattering angle. In Fig. 5(c), we emphasize the merging of two sheets of solutions along the rescattering ridges: the red and green sheets, in agreement with Figs. Fig. 6(a)(iii) and (iv) [lower arrow], as well as the yellow and black solution in agreement with Figs. 6(a)(i) and (ii)[upper arrow]. Finally, the arrow in Fig. 5(d) shows the point at which all four sheets merge.
@@
Figure 6: Final [panels(a)] and initial [panel (b)] momentum grid points reached by trajectories in the four branches of solutions which contribute to the rescattering ridge for a softened potential ($\alpha=1\times 10^{-2}$) and all other parameters as described in Fig. 2. The momentum regions associated with the trajectories $4_{s}^{(s)}$ $4_{s^{\prime}}^{(s)}$, $4_{l}^{(s)}$, $4_{l^{\prime}}^{(s)}$ are shown by yellow, green (short), black and red (long) scatter points, respectively. In panels (a) the red labels in the top right corners indicate the orbit types. The dashed lines at the boundaries of the scatter graphs indicate the merging of two or more sheets of solutions, and the colors indicate with which sheet the merging occur. In panel (a)(i), the upper (lower) dashed lines indicate a merging with the solution $4_{s^{\prime}}^{(s)}$ ($4_{l}^{(s)}$). In panel (a)(ii) the upper (lower) dashed lines indicates a merging with the solution $4_{l^{\prime}}^{(s)}$ ($4_{s}^{(s)}$). In panel (a)(iii) the upper (lower) dashed lines indicate a merging with the solution $4_{s}^{(s)}$ ($4_{l^{\prime}}l^{(s)}$). In panel (a)(iv) the upper (lower) dashed lines indicate a merging with the solution $4_{l}^{(s)}$ ($4_{s^{\prime}}l^{(s)}$). Panel (b) displays the initial momenta corresponding to the four sheets of solutions, with the additional solutions occupying the regions of very small perpendicular momenta $p_{0\perp}$. Both yellow and red points correspond to trajectories with an odd number of focal points while the black and green points correspond to trajectories with an even number of focal points.
@@
The initial momentum map, shown in Fig. 6(b) for the soft-core potential, is practically identical to its hard-core counterpart in the regions of large parallel momenta. However, close to the laser-polarization axis, there are significant differences, with the appearance of two more branches. This is expected from the discussion of Fig. 6(a). An interesting feature is that the red and green regions are much more localized in the $p_{0\parallel}p_{0\perp}$ plane than the black and yellow ones and approach them from below. This suggests a much stronger interaction with the core, with lower perpendicular momentum components. As the softening is increased, the region occupied by the green and red solutions widens. Furthermore, in contrast to what happens for the hard-core potential, they never touch the $p_{0\parallel}$ axis. These are precisely the momentum regions that would close the ridge in Figs. 6(a)(i)-(iv) and Fig. 3(b).
@@ IV Trajectory analysis and rescattering angles
Next, we will discuss why the softening of the potential influences the rescattering angle. We will start by looking at the electron trajectories in position and momentum space for the soft-core and the hard-core Coulomb potentials. These trajectories are displayed in Figs. 7(a) to (c), together with the Jacobian associated with each solution [Fig. 7(d)]. Although the hard-core potential exhibits only two branches of solutions, whose initial conditions are indicated in Figs. 4 and 6 by black and yellow scatter plots, in order to facilitate a comparison with the soft-core potential, in the figure we have slightly changed this convention and employed the same colors for orbits with the same initial conditions.
Figure 7: Position (panel (a)) and momentum (panel (b)) space trajectories and the inverse arcsinh of the Jacobian determinant along these trajectories (panel(d)) are plotted for a softened potential with $\alpha=1\times 10^{-2}$ (hard potential with $\alpha=1\times 10^{-6}$) using solid (dashed) lines with field parameters as described in Fig. 2. For the softened potential, the colors of the trajectories correspond to the color of the scatter point in Fig. 6 representing the boundary values of the trajectories momentum and the final momentum of each trajectory shown for the softened potential is $(p_{f\parallel},p_{f\perp})=(1,-1.2)$ atomic units. The initial momenta for the trajectories shown with the yellow, black, green and red lines are $(p_{0\parallel},p_{0\perp})=(-0.3949,0.01005)$, $(p_{0\parallel},p_{0\perp})=(0.09192,0.008766)$, $(p_{0\parallel},p_{0\perp})=(-0.3903,0.003019)$ and $(p_{0\parallel},p_{0\perp})=(0.09549,0.001135)$ atomic units, respectively for both the softened and hard potential. In panel (c), a contour plot of the difference between the soft and Coulomb potential (labeled as the Coulomb deviation) is shown with a more detailed view of some of the position space paths shown in panel (a) for which the same color and line-style scheme has been used. The parameters of the field are the same as those used in Fig. 2.
@@
For the hard-core potential, the trajectories shown have the same initial momentum as the trajectories in the softened potential, but the final momenta are all different. The differences between the $4_{s}^{(h)}$ ($4_{l}^{(h)}$) and the $4_{s}^{(s)}$ ($4_{l}^{(s)}$) trajectories with the same initial momenta are small but there is a significant difference for the $4_{s}^{(h)}$ ($4_{l}^{(h)}$) and the $4_{s^{\prime}}^{(s)}$ ($4_{l^{\prime}}^{(s)}$) trajectories. This is expected as the trajectories associated with the black and yellow branches in Fig. 6 do not deviate strongly from their Coulomb-type counterparts. For the softened potential, although the position space trajectories in Fig. 7(a) begin on very different paths, eventually the red (green) trajectory coincides precisely with the black (yellow) trajectory. The pairwise convergence of trajectories in different branches to the same asymptotic coordinates is a consequence of their long-distance behavior only having two possible branches of solutions, as it is determined by the Coulomb tail.
The non-Coulomb influence is also visible in momentum space, displayed in Fig. 7(b). The black and the yellow soft-core trajectories are nearly circular, and behave like their Coulomb counterparts. In contrast, the green and red trajectories are elliptical for the soft-core potential and nearly circular for the Coulomb potential. The circular shape observed for the Coulomb orbits stems from the conserved quantity associated with the dynamical symmetry of the Coulomb potential, which is the Laplace-Runge-Lenz (LRL) vector. In the pure Coulomb dynamics, LRL vector conservation implies that paths in momentum space are circular and they have a radius given by $1/L$, where $L$ is the angular momentum, and a center defined by the LRL vector [57, 58, 59]. Because rescattering occurs around a field zero crossing, this symmetry is approximately observed for scenarios in which the Coulomb-type dynamics are prevalent. This renders the momentum-space orbit approximately circular and happens if, during the rescattering process, the trajectory remains at a distance from the core greater than the length scale comparable with the softening parameter $\alpha$. Examples are the yellow and black orbits, which remain very similar and nearly circular for both potentials. However, this symmetry is unambiguously broken once the non-Coulomb dynamics determined by the softening becomes more important. This is observed for the green and red orbits, which reach closer to the core than $\alpha$, and whose momentum paths vary dramatically from being circular.
@@
Finally, in Fig. 7(d) we present the Jacobian determinant as a function of time for specific orbits around the instance of rescattering. This is a very short time interval but it contains a large amount of the electrons’ dynamics. For trajectories that pass close enough to the core to be altered significantly by the absence of the Coulomb singularity (green and red orbits), there is a large spike in $\textrm{det}(J)$ towards $\infty$- ∞ followed by a spike towards +$\infty$+ ∞. Conversely, for trajectories that do not pass as close to the core, so that they are similar to those of electrons in an actual Coulomb potential (black and yellow orbits), there is just a single large spike in $\textrm{det}(J)$ towards $\infty$- ∞.
@@
The role of the softening in the potential can be further explored using scattering theory. Figs. 8(a) and (b) provide a schematic representation of a scattering process by a hard-core ($\alpha=10^{-6}$) and a soft-core ($\alpha=10^{-2}$) Coulomb potential. The figures show that, for a hard-core Coulomb potential, an incoming electron can be rescattered up to an angle close to $\pi$, while, for a soft-core potential, the particle will not feel the potential and will pass through it for angles larger than a certain critical value. Explicit expressions for the scattering angles of a Coulomb and soft-core potential have been provided in [60] for the field-free case. A schematic representation of the scattering process, as well as the relevant parameters, is shown in Fig. 9. In the center of mass frame, the scattering angle $\phi$ can be expressed as,
@@
In Fig. 8(c), we plot the scattering angle as a function of the impact parameter for initial kinetic energy corresponding to the $10.9U_{p}$ ridge. For both the hard-core and the soft-core Coulomb potential, the angle exhibits a maximum and then decays as the impact parameter approaches zero. The maximum is sharper and closer to $\pi$ for the hard-core Coulomb potential. As the softening increases, the maximum occurs at a larger impact parameter, and as the impact parameter is reduced further, the scattering angle decreases. In the limit of vanishing softening, as we approach the exact Coulomb potential, the angle decreases monotonously as the impact parameter increases, reaching a maximum value of $\pi$ as the impact parameter approaches zero [shown with the dotted brown line in Fig. 8(c)]. This sheds light on the behavior observed in the upper panels of the figure. For the soft-core potential, as the impact parameter approaches zero, the trajectories pass by without feeling the potential.
In Fig. 8(d), the dots represent the maximum angle $\theta$ defined by the red arrows in Fig. 2 and represented in the schematic model in Fig. 9, for the three shortest pairs of backscattered trajectories as a function of the initial kinetic energy for each value of the softening parameter. For the hard-core Coulomb potential, we have set the value to zero. With lines, we plot the results obtained from our scattering model, where $\theta$ can be approximately defined in terms of the scattering angle $\phi$, the peak of the vector potential $\mathbf{A}_{0}$ and the scattering kinetic energy $k_{e}$ as
@@
The predictions of the scattering model follow the same trends as the angle estimated from the PMDs in Fig. 2, although the agreement is better for the first return leading to the high-energy ridge. Intuitively this could be explained if we consider that the trajectories contributing to the first return are closer to the scattering model, while for later returns the influence of the residual potential for a longer time prior to rescattering precludes determining a clear $t_{\mathrm{resc}}$. This issue also implies that Born-type theories such as the SFA exhibit a worse agreement with fully Coulomb-distorted approaches for longer orbits [32]. The values obtained agree reasonably well with the white arrows in Fig. 2($c^{\prime}$).
Figure 9: A schematic showing the simplified model used to understand the behavior of trajectories in the presence of a softened potential. It shows the process of converting from the scattering angle $\phi$ computed from the field-free spherically symmetric scattering problem, to the angle $\theta$ which we observe in our photoelectron momentum distributions. The inset panel shows the type of shift which occurs to the rescattering ridges by approximately $\mathbf{A}_{0}$ due to the field dressing of the scattered electron.

### v2

@@ III.2 Momentum mapping and catastrophes
For simplicity, in the figures that follow, we will focus on the orbits leading to the primary rescattering ridge whose energy is around $10U_{p}$, but the arguments employed in the subsequent analysis are general and can be applied to the remaining pairs of orbits. Furthermore, we have considered the initial momenta in the upper half-plane. For the orbits associated with the primary rescattering ridge, this leads to final momenta in the lower half plane, of which we just show a single quadrant (see upper panels in Fig. 5, Fig. 6 and upper panels in Fig. 7). In Fig. 3, we have used the upper half plane for the final momenta to construct the PMD, which means that the initial momenta arein the lower half-plane. However, this does not alter the subsequent conclusions as the PMDs are symmetric with regard to a reflection upon the $p_{f\parallel}$ axis. For the specific case of a monochromatic field, we also expect a reflection symmetry with regard to the $p_{f\perp}$ axis for the momentum mapping. Therefore, we have only included ionization times in a single half-cycle without loss of generality.
@@
We have observed that, for this particular pair of orbits and potential, the short (long) orbits have an odd (even) number of focal points, each of which occurs at a time when the Jacobian is vanishing. Although we cannot yet make concrete statements about the physics, an electron returning before or after a field zero crossing may influence its dynamics in relation to focal points. That being said, the existence of the fold catastrophe between the respective sheets of solutions containing the long and short trajectories, does require the existence of an additional focal point in one of them.
It is worth noting that not all focal points will cause relevant dynamical changes in the system. Some focal points occur in pairs, so that the Maslov phases associated with them will cancel each other. They are not directly related to rescattering. However, the remaining focal points will appear at the same time as a spike in the Jacobian, which is associated with rescatering and is important for the dynamics. Both orbits undergo one act of rescattering, and exhibit a single spike in the Jacobian. Physically, a spike indicates sudden changes in the electron momentum. A summary of these findings is provided in Table 2.
The corresponding initial momentum maps for the hard-core potential are given in Fig. 5(b), where the same color code as for the final momenta was used. The overall momenta for the long orbit is much smaller, which is consistent with ionization times being closer to the peak of the field. Moreover, the density of points is highest close to the polarization axis and is lowest near and at the boundary between the black and yellow regions. The high density is in line with the expected dynamics for a linearly polarized field, as ionization will happen predominantly close to the axis. Nonetheless, these regions are distorted by the presence of the Coulomb potential.
@@
Figs. 7(a)(i)-(iv) show the final momentum maps associated with each of these solutions. A noteworthy feature is that there are no solutions close to the polarization axis. This confirms that the behavior observed for the PMD in Fig. 3 is due to a change in kinematic constraints. The figure also shows that, for the soft-core case, different sheets of solutions obtained for the final momenta merge in several places, not only the rescattering ridge. The boundaries at which different solutions are glued together are represented by the dashed lines in each panel, and the color indicates to which solution it is connected. For instance, in Fig. 7(a)(i), we see that the solution $4^{(s)}_{s}$, whose final momenta are represented by the yellow scatter plot, still merges at the ridge with the long orbit $4^{(s)}_{l}$, whose final momentum mapping is given by the black scatter plot in Fig. 7(a)(ii). This behavior resembles that of their hard-core counterparts. However, in addition to that, $4^{(s)}_{s}$ is also connected to the solution $4^{(s)}_{s^{\prime}}$ [Fig. 7(a)(iii)] as the green dashed line indicates. The solution $4^{(s)}_{s^{\prime}}$, on its turn, is also connected to the $4^{(s)}_{l^{\prime}}$ solution [shown in Fig. 7(a)(iv)] via the rescattering ridge. Solutions $4^{(s)}_{l^{\prime}}$ and $4^{(s)}_{l}$ coalesce at the boundary defined near the minimal rescattering angle. All four solutions merge at a single point, where the four folds which characterise the surface of solutions intersect..
@@
Figure 6: Three dimensional representation of the initial perpendicular momentum map as a function of the electron’s final momentum components, calculated for the same parameters as in the previous figures, and a softening of $\alpha=1\times 10^{-6}$ [panel (a)] and $\alpha=1\times 10^{-2}$ [panels(b), (c) and (d)]. The colors of the surfaces match those used in the two-dimensional projections in Figs. 5 and 7, and the arrows indicate the merging of two or more sheets.
@@
This merging is better seen in the full three-dimensional surface, displayed in Figs. 6(b) to (d). The arrows in Fig. 6(b) indicate the coalescence of two solutions along the fold introduced by the softening. The left arrow shows the merging of the red and black branches of solutions, closest to the $p_{f,\parallel}$ axes, while the right arrow indicates the fusion of the yellow and green branches. The coalescence of these sheets is determined by the minimal rescattering angle. In Fig. 6(c), we emphasize the merging of two sheets of solutions along the rescattering ridges: the red and green sheets, in agreement with Figs. Fig. 7(a)(iii) and (iv) [lower arrow], as well as the yellow and black solution in agreement with Figs. 7(a)(i) and (ii)[upper arrow]. Finally, the arrow in Fig. 6(d) shows the point at which all four sheets merge.
@@
Figure 7: Final [panels(a)] and initial [panel (b)] momentum grid points reached by trajectories in the four branches of solutions which contribute to the rescattering ridge for a softened potential ($\alpha=1\times 10^{-2}$) and all other parameters as described in Fig. 2. The momentum regions associated with the trajectories $4_{s}^{(s)}$ $4_{s^{\prime}}^{(s)}$, $4_{l}^{(s)}$, $4_{l^{\prime}}^{(s)}$ are shown by yellow, green (short), black and red (long) scatter points, respectively. In panels (a) the red labels in the top right corners indicate the orbit types. The dashed lines at the boundaries of the scatter graphs indicate the merging of two or more sheets of solutions, and the colors indicate with which sheet the merging occurs. In panel (a)(i), the upper (lower) dashed lines indicate a merging with the solution $4_{s^{\prime}}^{(s)}$ ($4_{l}^{(s)}$). In panel (a)(ii) the upper (lower) dashed lines indicates a merging with the solution $4_{l^{\prime}}^{(s)}$ ($4_{s}^{(s)}$). In panel (a)(iii) the upper (lower) dashed lines indicate a merging with the solution $4_{s}^{(s)}$ ($4_{l^{\prime}}^{(s)}$). In panel (a)(iv) the upper (lower) dashed lines indicate a merging with the solution $4_{l}^{(s)}$ ($4_{s^{\prime}}^{(s)}$). Panel (b) displays the initial momenta corresponding to the four sheets of solutions, with the additional solutions occupying the regions of very small perpendicular momenta $p_{0\perp}$. Both yellow and red points correspond to trajectories with an odd number of focal points while the black and green points correspond to trajectories with an even number of focal points.
@@
The initial momentum map, shown in Fig. 7(b) for the soft-core potential, is practically identical to its hard-core counterpart in the regions of large perpendicular momenta. However, close to the laser-polarization axis, there are significant differences, with the appearance of two more branches. This is expected from the discussion of Fig. 7(a). An interesting feature is that the red and green regions are much more localized in the $p_{0\parallel}p_{0\perp}$ plane than the black and yellow ones and approach them from below. This suggests a much stronger interaction with the core, with lower perpendicular momentum components. As the softening is increased, the region occupied by the green and red solutions widens. Furthermore, in contrast to what happens for the hard-core potential, they never touch the $p_{0\parallel}$ axis. These are precisely the momentum regions that would close the ridge in Figs. 7(a)(i)-(iv) and Fig. 3(b).
@@ IV Trajectory analysis and rescattering angles
Next, we will discuss why the softening of the potential influences the rescattering angle. We will start by looking at the electron trajectories in position and momentum space for the soft-core and the hard-core Coulomb potentials. These trajectories are displayed in Figs. 8(a) to (c), together with the Jacobian associated with each solution [Fig. 8(d)]. Although the hard-core potential exhibits only two branches of solutions, whose initial conditions are indicated in Figs. 5 and 7 by black and yellow scatter plots, in order to facilitate a comparison with the soft-core potential, in the figure we have changed this convention and employed the same colors for orbits with the same initial conditions. However, physically, the red (green) dashed orbits associated with the hard-core potential in Fig. 8 leave from the black (yellow) momentum regions in Fig. 5(b).
Figure 8: Position (panel (a)) and momentum (panel (b)) space trajectories and the inverse arcsinh of the Jacobian determinant along these trajectories (panel(d)) are plotted for a softened potential with $\alpha=1\times 10^{-2}$ (hard potential with $\alpha=1\times 10^{-6}$) using solid (dashed) lines with field parameters as described in Fig. 2. For the softened potential, the colors of the trajectories correspond to the color of the scatter point in Fig. 7 representing the boundary values of the trajectories momentum and the final momentum of each trajectory shown for the softened potential is $(p_{f\parallel},p_{f\perp})=(1,-1.2)$ atomic units. For the hard-core potential, we have used the same color codes for trajectories with the same initial conditions to facilitate direct comparison with the soft-core counterparts, although, physically, the initial conditions of the green and red dashed trajectories occupy the yellow and black regions of the mapping in Fig. 5(b). The same convention has been used for the Jacobian determinants in panel (d) of the present figure. The initial momenta for the trajectories shown with the yellow, black, green and red lines are $(p_{0\parallel},p_{0\perp})=(-0.3949,0.01005)$, $(p_{0\parallel},p_{0\perp})=(0.09192,0.008766)$, $(p_{0\parallel},p_{0\perp})=(-0.3903,0.003019)$ and $(p_{0\parallel},p_{0\perp})=(0.09549,0.001135)$ atomic units, respectively for both the softened and hard potential. For the black and yellow trajectories used in this figure, there is a good similarity between the trajectories from the soft and hard potentials, which is why they overlap and it appears as if there are only two trajectories plotted for the hard potential. In panel (c), a contour plot of the difference between the soft and Coulomb potential (labeled as the Coulomb deviation) is shown with a more detailed view of some of the position space paths shown in panel (a) for which the same color and line-style scheme has been used. The parameters of the field are the same as those used in Fig. 2.
@@
For the hard-core potential, the trajectories shown have the same initial momentum as the trajectories in the softened potential, but the final momenta are all different. The differences between the $4_{s}^{(h)}$ and the $4_{s}^{(s)}$ trajectories, represented by the yellow dashed and solid lines, respectively, are minimal, so that the curves overlap in the figure. The same holds for the $4_{l}^{(h)}$ and $4_{l}^{(s)}$ trajectories with the same initial momenta, which are represented by black lines. This is expected as the trajectories associated with the black and yellow branches in Fig. 7 do not deviate strongly from their Coulomb-type counterparts. However, there is a significant difference for the $4_{s}^{(h)}$ ($4_{l}^{(h)}$) and the $4_{s^{\prime}}^{(s)}$ ($4_{l^{\prime}}^{(s)}$) trajectories [red and green lines]. For the softened potential, although the position space trajectories in Fig. 8(a) begin on very different paths, eventually the red (green) trajectory coincides precisely with the black (yellow) trajectory. The pairwise convergence of trajectories in different branches to the same asymptotic coordinates is a consequence of their long-distance behavior only having two possible branches of solutions, as it is determined by the Coulomb tail.
The non-Coulomb influence is also visible in momentum space, displayed in Fig. 8(b). The black and the yellow soft-core trajectories are nearly circular, and behave like their Coulomb counterparts. In contrast, the green and red trajectories are elliptical for the soft-core potential and nearly circular for the Coulomb potential. The circular shape observed for the Coulomb orbits stems from the conserved quantity associated with the dynamical symmetry of the Coulomb potential, which is the Laplace-Runge-Lenz (LRL) vector. In the pure Coulomb dynamics, LRL vector conservation implies that paths in momentum space are circular and they have a radius given by $1/L$, where $L$ is the angular momentum, and a center defined by the LRL vector [57, 58, 59]. Because rescattering occurs around a field zero crossing, this symmetry is approximately observed for scenarios in which the Coulomb-type dynamics are prevalent. This renders the momentum-space orbit approximately circular and happens if, during the rescattering process, the trajectory remains at a distance from the core greater than the length scale comparable with the softening parameter $\alpha$. Examples are the yellow and black orbits, which remain very similar and nearly circular for both potentials. However, this symmetry is unambiguously broken once the non-Coulomb dynamics determined by the softening become more important. This is observed for the green and red orbits, which reach closer to the core than $\alpha$, and whose momentum paths vary dramatically from being circular.
@@
Finally, in Fig. 8(d) we present the Jacobian determinant as a function of time for specific orbits around the instance of rescattering. This is a very short time interval but it contains a large amount of the electrons’ dynamics. For trajectories that pass close enough to the core to be altered significantly by the absence of the Coulomb singularity (green and red orbits), there is a large spike in $\textrm{det}(J)$ towards $\infty$- ∞ followed by a spike towards +$\infty$+ ∞. Conversely, for trajectories that do not pass as close to the core, so that they are similar to those of electrons in an actual Coulomb potential (black and yellow orbits), there is just a single large spike in $\textrm{det}(J)$ towards $\infty$- ∞. The first focal point is common to both hard- and soft-core potentials and is associated with the momentum component perpendicular to the driving-field polarization changing sign due to the residual binding potential. This occurs at a length scale in which the soft- and hard-core potentials are virtually indistinguishable. For trajectories which pass closest to the minimum of the soft-core potential, there exists an additional focal point. While for the hard-core potential the trajectories follow the standard hyperbola associated with Coulomb-type behavior, for the soft-core potential the attracting force is weaker, especially close to the core. Therefore, there will be less deflection, some of the trajectories will disperse as seen in Fig. 8(c) and defocusing will occur.
@@
The role of the softening in the potential can be further explored using scattering theory. Figs. 9(a) and (b) show the field-free trajectories for a hard-core ($\alpha=10^{-6}$) and a soft-core ($\alpha=10^{-2}$) Coulomb potential, respectively, as a function of the impact parameter, indicated by the color bar in the figure. The figures show that, for a hard-core Coulomb potential, as the impact parameter decreases, an incoming electron can be rescattered up to an angle close to $\pi$. However, for the soft-core potential, as the impact parameter approaches zero, the trajectories pass through the core without deflection.
Explicit expressions for the scattering angles of a Coulomb and soft-core potential have been provided in [60] for the field-free case. A schematic representation of the scattering process, as well as the relevant parameters, is shown in Fig. 10. In the center of mass frame, the scattering angle $\phi$ can be expressed as,
@@
In Fig. 9(c), we plot the scattering angle as a function of the impact parameter for initial kinetic energy corresponding to the $10.9U_{p}$ ridge. For both the hard-core and the soft-core Coulomb potential, the angle exhibits a maximum and then decays as the impact parameter approaches zero. The maximum is sharper and closer to $\pi$ for the hard-core Coulomb potential. As the softening increases, the maximum occurs at a larger impact parameter, and as the impact parameter is reduced further, the scattering angle decreases. In the limit of vanishing softening, as we approach the exact Coulomb potential, the angle decreases monotonously as the impact parameter increases, reaching a maximum value of $\pi$ as the impact parameter approaches zero [shown with the dotted brown line in Fig. 9(c)]. This sheds light on the behavior observed in Fig 9(b). For the soft-core potential, as the impact parameter approaches zero, the scattering angle goes to $\phi=\pi/2$ [see solid line in Fig. 9(c)]. From the schematic representation in Fig. 10, this corresponds to a straight line, as the ones represented in Fig. 9(b).
In Fig. 9(d), the dots represent the minimum angle $\theta$ defined by the red arrows in Fig. 3 and represented in the schematic model in Fig. 10, for the three shortest pairs of backscattered trajectories as a function of the initial kinetic energy for each value of the softening parameter. For the hard-core Coulomb potential, we have set the value to zero. With lines, we plot the results obtained from our scattering model, where $\theta$ can be approximately defined in terms of the scattering angle $\phi$, the peak of the vector potential $\mathbf{A}_{0}$ and the scattering kinetic energy $k_{e}$ as
@@
The predictions of the scattering model follow the same trends as the angle estimated from the PMDs in Fig. 3, although the agreement is better for the first return leading to the high-energy ridge. Intuitively this could be explained if we consider that the trajectories contributing to the first return are closer to the scattering model, while for later returns the influence of the residual potential for a longer time prior to rescattering precludes determining a clear $t_{\mathrm{resc}}$. This issue also implies that Born-type theories such as the SFA exhibit a worse agreement with fully Coulomb-distorted approaches for longer orbits [32]. The values obtained agree reasonably well with the white arrows in Figs. 3((b)-3(d).
Figure 10: A schematic showing the simplified model used to understand the behavior of trajectories in the presence of a softened potential. It shows the process of converting from the scattering angle $\phi$ computed from the field-free spherically symmetric scattering problem, to the angle $\theta$ which we observe in our photoelectron momentum distributions. The inset panel shows the type of shift that occurs to the rescattering ridges by approximately $\mathbf{A}_{0}$ due to the field dressing of the scattered electron.
Fig. 11 presents the outcome of the analytical model for the minimal angle (lines) calculated for the outermost scattering ridge, plotted against the results read directly from the PMDs (scatter plots). Besides the very good agreement, the results also show that the minimal angle $\theta$ increases with the driving-field wavelength. This at first sight surprising result is discussed in more detail in the Appendix A, and is caused by longer wavelengths concentrating the initial momentum distributions close to the polarization axis. This leads to a stronger interaction with the core upon rescattering, so that the non-Coulomb effects become more relevant.
Figure 11: Minimal $\theta$ determined by the scattering model outlined in this section is plotted against the wavelength, $\lambda$, of the driving field, for the outermost rescattering ridge. The different lines represent different values of the photoelectron energy in units of $U_{P}$ which itself depends on $\lambda$. The maximum photoelectron energy $E^{(1)}_{ridge}$ for the shortest pair of orbits in units of $U_{p}$ has a $\lambda$-dependence. For $\lambda=2000$ nm, $E^{(1)}_{ridge}=10.13U_{p}$, for $\lambda=1300$ nm, $E^{(1)}_{ridge}=10.34U_{p}$ and for $\lambda=800$ nm, $E^{(1)}_{ridge}=10.86U_{p}$ (see Table I). Lines are plotted for the maximum photoelectron energy corresponding to each of the values of $\lambda$ for which PMDs were computed using the CQSFA. The PMDs for $\lambda=1300$ nm and 2000 nm are shown in the appendix. The scatter points correspond to the angle $\theta$ calculated directly from the PMDs. The shaded areas indicate the angular range associated with the maximum photoelectron energy (around $10.13U_{p}\leq E^{(1)}_{ridge}\leq 10.86U_{p}$) for the outermost ridge in the range of wavelengths considered in the figure.